OpenStudy (allymccoy28):

Solve for x: the quantity of x plus 20 over 2 = 3x

OpenStudy (lochana):

\[\frac{x+20}{2}\]

OpenStudy (jazzyluvsu123):

x+202=3x 2(x+202)=3x(2) This cancels out the 2 in the fraction x+20=6x -x -x 20 = 5x /5 /5 x = 4

OpenStudy (jazzyluvsu123):

x+20=3x

OpenStudy (jazzyluvsu123):

\[2(x+20)=3x(2) This cancels out the 2 \in the fraction x+20=6x -x -x 20 = 5x /5 /5 x = 4\]

OpenStudy (jazzyluvsu123):

Here - x + 20/2 = 3x multiple both sides by 2 to get rid of the 2 2(x+20/2) = 3x(2) the 2s cancel each other out on the left side x+20=6x you want to get the x's on the same side so subtract x from each side 20=5x divide both sides by 5 to get x on its own 20/5 = 5x/5 4=x

OpenStudy (lochana):

mm.. I thought it is (x+20)/2